Jon Paul Nollmann

né Darren Senn Ballistic Technologies sinster@balltech.net (preferred) (888) 869-5668 (pager)

(510) 719-9963 (cell)

Education

Bachelor’s Degree, Computer Engineering. Honors in the Comprehensive. UCSC. Self educated in nearly all my skills. Past projects

16-bit fully orthogonol CPU (discrete)

UART (integrated logic)

2 1/2 octave 2 voice programmable wavetable organ with adjustable attack, decay, and mixing. (discrete

logic) Handheld automotive diagnostics computer. Database replication daemon.

Employment

5/96-Present: Ballistic Technologies

Running my own business to do contract hardware and software engineering. Clients have included: SCO: Network appliance security audit. AdAce: linux system architecting, security design, credic card processing system development, and

database security. Siebel Systems: Call center application maintenance and improvements. SQL optimization. NT and

UNIX portable C++, SQL, Install Shield, TCL. Energy Interactive: Firewall, proxy server, and mail host configuration. W3 Corporation: Linux system installation, repair, internal security audit, and system administration

consultation. Predictive Systems, Inc: Security software installation and integration on NT, Solaris, and HP/UX. Japan Systems America: Linux system repair and system administration consultation. Scopus International: Motif/X/UNIX debugging on IRIX, SunOS, Solaris, and HP/UX. Internationalization. Database programming. Development of an offline database replication daemon under NT and UNIX. System security consultation.

Interro Systems: Instruction of engineers on real time programming with a specific emphasis on Interro’s

product. See below. Thuridion: Motif/X/UNIX debugging on Solaris. Postscript debugging. Internationalization.

6/92-5/96: Interro Systems, Inc.

One of the original engineers, I was central in the development of their entire product line, a family of handheld automotive diagnostic computers. I developed all the software which drives the acquisition circuitry, the data analysis code, and some of the user interface. I also contributed to the design of the acquisition hardware itself, including the period counter, gain circuitry, backlight, power monitoring, high speed acquisition, peak detect, and RS-232 communications. Configured and administered Interro’s computer resources, including their Internet connection, gateway, and web site.

10/89-6/92: University of California, Santa Cruz. User interface programmer for a research project developing an automated surface router for multichip modules, called SURF. X, Motif, and Postscript under various UNIX platforms.

10/87-10/89: University of California, Santa Cruz Student computer consultant. Ran one of the computer labs on campus, helping students, faculty and staff to use the systems. Taught courses ranging from the basic use of DOS to programming under UNICOS on the Cray Y/MP.

Hardware projects

POCSAG decoder Cell phone jammer Robotic controller Digital puzzlebox Yaesu FT-50R reprogrammer

High energy broadband RF cannon

Software projects

Linux (one of the original kernel authors) Yaesu FT-50R enhancements

Text projects

Darkwater GIT Lessons (http://www.darkwater.com/git/index.html)

Other projects

Sysadmin for Darkwater and my own machines for the last 10 years, and webmaster for Darkwater for the past 3 years, helping with the administration of dis.org for the past year and a half. Former domain registrar for all the .ca.us domains within Santa Cruz and southern Santa Clara counties.

Hobbies

Hunting, target shooting, pyrotechnics, chemistry, electronics, fencing, video games, cooking (and eating!), reading, cryptography, legal wrangling, motorcycles, swimming.

Resume of JP Nollmann. Page 2